# Street Gaussians for Modeling Dynamic Urban Scenes

### v1

@@ 4 Implementation details
Initialization.  We use LiDAR point cloud captured by ego vehicle as initialization. Specifically, we first collect points inside the bounding boxes and transform them into the local coordinate system to initialize each object model. Then we downsample the remaining aggregated point cloud with a voxel size of 0.15m and filter out invisible points for the background model. The colors of LiDAR point cloud are obtained by projecting to the corresponding image plane and querying the pixel value. Due to the limited coverage of LiDAR point clouds over large areas, we also incorporate SfM [39] point cloud to handle distant regions.
Densification and pruning.  We follow [16] to apply adaptive control during optimization. Original 3D Gaussians set the scene radius based on camera poses, which is not suitable for large-scale urban scenes. As a result, we obtain the radius of the background model by the LiDAR point cloud while the radius of the object model is determined by the bounding box scale. In order to prevent object Gaussians from growing to occluded areas, for each object model we sample a set of points as a probability distribution function. During optimization, Gaussians with sampled points outside the bounding box will be pruned.
@@ 5.1 Experimental Setup
We train Street Gaussians for 30000 iterations with Adam optimizers [17] following the learning rate configurations of 3D Gaussians [16]. The learning rate of translation transformation $\Delta\mathbf{T}_{t}$ and rotation transformation $\Delta\mathbf{R}_{t}$ are set to 0.005 and 0.001, which decay exponentially to $5e^{-5}$ and $1e^{-5}$ respectively. All the experiments are conducted on one RTX 4090 GPU.
Datasets.  We conduct experiments on Waymo Open Dataset [44] and KITTI benchmarks [13]. On the Waymo Open Dataset, we select 6 recording sequences with large amounts of moving objects, significant ego-car motion and complex lighting conditions. All sequences have a length of around 100 frames. We select every 10th image in the sequence as the test frames and use the remaining for training. As we find that our baseline methods [33, 55] suffer from high memory cost when training with high-resolution images, we downscale the input images to $1066\times 1600$. On KITTI [13] and Vitural KITTI 2 [6], we follow the settings of MARS [55] and evaluate our methods with different train/test split settings. We use the bounding boxes generated by the detector[54] and tracker [53] on Waymo dataset and use the officially provided object tracks from KITTI.
Baseline methods.  We compare our methods with three recent methods. (1) NSG [33] represents background as multi-plane images and use per-object learned latent codes with a shared decoder to model moving objects. (2) MARS [55] builds the scene graph based on Nerfstudio [46]. (3) 3D Gaussians [16] models the scene with a set of anisotropy gaussians. Both NSG and MARS are trained and evaluated using ground truth bounding boxes, we try different versions of their implementations and report the best result for each sequence. We also replace the SfM point cloud in 3D Gaussians with the same input as our method for fair comparison. See supplementary material for details.
Figure 4: Qualitative comparisons on the Waymo [44] dataset. NSG [33] and MARS [55] often produce blurry and distorted results. 3D GS [16] fails to handle dynamic objects and generates ghosting artifacts in these cases. In contrast, our approach significantly outperforms other methods with high fidelity and sharp details.
Figure 5: Qualitative comparisons on the KITTI [13] dataset.
3D GS [16] NSG [33] MARS [55] Ours
[…]
PSNR*$\uparrow$ 17.74 22.05 23.07 25.16
@@
5.2 Comparisons with the State-of-the-art
Tables 1, 2 present the comparison results of our method with baseline methods [16, 33, 55] in terms of rendering quality and rendering speed. We adopt PSNR, SSIM and LPIPS [68] as metrics to evaluate rendering quality. To better evaluate the quality of moving objects, we project 3D bounding boxes to 2D image plane and calculate the loss only on pixels inside the projected box, which is denoted as PSNR* in our experiments. For all the metrics, our model achieves the best performance among all the methods. Moreover, our method renders two magnitudes faster than MARS and NSG. Although 3D GS is faster than our method, it can only support static scenes and the result of moving objects degrades significantly.
Figures 4, 5 show the qualitative results of our method and baselines. NSG and MARS suffer from blurry and distorted results, which is especially marked when the camera has a large motion. This should be attributed to their model lacking sufficient expressive capability when the scene is complex. 3D GS fails to handle dynamic objects and generates ghosting artifacts in these cases. In contrast, our method can generate high-quality novel views with high fidelity and details.
@@ 5.2 Comparisons with the State-of-the-art
Figure 6: Visual ablation results on the Waymo [44] dataset. The results indicate that optimizing tracked poses improves the quality. “opt.” denotes optimization.
@@ 5.3 Ablations and Analysis
We validate our algorithm’s design choices on all six select sequences from the Waymo dataset. Tables 3 present the quantitative results. The influence of optimizing tracked poses.  Experimental results in Table 3 show that without optimizable tracked poses the rendering quality degrades much. This is because inaccurate poses make the object model supervised under the wrong pixel value. Figure 3 shows some visual results. To further validate the effectiveness of optimizing tracked poses, we compare our method with a variant that uses ground truth poses from the dataset. As shown in row2 Table 3, the performance gap between our complete model and the model trained with ground truth poses is small, which indicates the effectiveness of our pose optimization strategy. It is interesting to notice that the LPIPS metric of our complete model issue even better than the model trained with ground truth poses, a plausible explanation is that tracked poses of real-world data is obtained from manual annotations where noisy results still exist.
[…]
Ours w/o pose opt. 34.05 23.37 0.934 0.094
Ours w/o 4DSH 34.38 24.72 0.935 0.093
Ours w/ GT pose 34.57 25.24 0.936 0.092
Complete model 34.54 25.16 0.936 0.091
Visual results of the influence of tracked pose optimization is shown in Figure 6. The first two rows indicate that treating tracked poses as learnable parameters help the object model synthesize more texture details and reduce rendering artifacts. Models trained directly with tracked poses suffer from inaccurate positions in the world frame. Moreover, In comparison with images rendered with ground truth poses, our results with optimized tracked poses are comparable or even better on challenging regions like the rear of the white vehicle or the logo of the black vehicle in Figure 6.
@@
Figure 7: Decomposition results on the KITTI [13] dataset. NSG [33] cannot decompose clean foreground objects while PNF [19] produces blurry results. Instead, our method successfully decomposes the foreground objects and produces high fidelity rendering results.
5.4 Applications
Figure 8: Editing operations on the Waymo [44] dataset. Images in the left and right columns represent the results before and after editing. Our method supports various editing operations, including translation, rotation and swapping.
Street Gaussians can be applied to multiple tasks in computer vision including object decomposition, semantic segmentation and scene editing. Object Decomposition.  We compare the decomposition results of our method with NSG [33] and PNF [19] under KITTI dataset. As shown in Figure 7, NSG fails to disentangle foreground objects from the background and the result of PNF is lack of details. In contrast, our method can produce high-fidelity decomposed rendering results.
Scene editing.  Our scene representation enables various types of scene editing operations. We can translate the vehicle (Figure 8 (a)), rotate the heading of the vehicle (Figure 8 (b)) and swap the vehicles with another one (Figure 8 (c)).
Semantic Segmentation.  We compare the quality of our rendered semantic map with the semantic prediction from Video-K-Net [20] on KITTI dataset. Our semantic segmentation model is trained with results from Video K-Net. Qualitative and quantitative results are shown in Figure 9 and Table 4. Our rendered semantic maps achieve better performance thanks to our 3D semantic representation.
Figure 9: Visual semantic segmentation results on the KITTI [13] dataset. It can be observed that our method achieves better performance, particularly in ambiguous areas such as shadows due to our ability to fuse semantic information in 3D.
@@ 5.4 Applications
6 Conclusion and Discussion
@@
This work also has some known limitations. First, our method is limited to reconstructing rigid dynamic scenes, such as static streets with only moving vehicles, and cannot handle non-rigid dynamic objects like walking pedestrians. Future work could consider employing more complex dynamic scene modeling methods [51], to address this issue. Second, the proposed method is dependent on the recall rate of off-the-shelf trackers. If some vehicles are missed, our pose optimization strategy cannot compensate for this. We leave this problem to future work.
Appendix
Appendix A More implementation details
Baselines.
On Waymo Open Dataset [44], we find that the results of NSG [33] and MARS [55] is better when using a separate neural field for each object compared with shared decoder and per-object optimized latent codes, especially when the number of moving objects is relatively high. One plausible reason is that the tracking results provided by the Waymo dataset do not differentiate between different types of vehicles, Therefore, the shared decoder needs to fit objects with significant differences in shape, which makes the optimization process more complex. As a result, We use the original NeRF [30] and Nerfacto [46] to represent each object model in NSG and MARS respectively. For MARS, we also add depth loss following their implementation.
Hyperparameters.
We set the number of fourier coefficients $k$ as 5 to maintain a balance between performance and storage cost. Due to the relatively less intense view-dependent effect on urban scene compared to dataset in 3D Gaussian [16], we reduce the SH degree to 1 to prevent overfitting. For the remaining parameters, we use the default values from the official implementation of 3D Gaussian [16].
Evaluation metrics.
Figure 10 visually illustrates the calculation method of the PSNR* metric in our experiments. For fair comparison, both our method and the baselines are evaluated using the mask obtained from ground-truth tracked boxes.
Regularization loss.
The regularization term in our loss function is defined as an entropy loss on the accumulated alpha values of decomposed foreground objects which can be generated in parallel during training. We add this loss after the adaptive control process to help Street Gaussians better distinguish foreground and background following [43]. Figure 11 shows the qualitative results, which demonstrates the effect of this regularization term.
Object semantic.
@@ Appendix A More implementation details
We learn a confidence score between 0 and 1 for each Gaussian in the object model. The logits of the vehicle semantic class output by each Gaussian is a constant $C$ multiplied by this confidence score. We set $C$ as 100 in our experiments. To merge the one-dimensional scalar $\beta_{o}$ with the M-dimensional vector $\beta_{b}$ of background, we convert $\beta_{o}$ to a M-dimensional one-hot vector for the vehicle label during rendering.
@@
Appendix B Additional experiments
Influence of initial point cloud.
We train one sequence from Waymo dataset using LiDAR point cloud or SfM point cloud as initialization respectively. The SfM point cloud of background is built by running Colmap [39] on input training images. We randomly sample points inside the bounding box for moving objects as we find that the recovered point cloud from Colmap is very sparse.
Figure 12 and Table 5 show the qualitative and quantitative results. Our method can generate high-fidelity images using either methods. SfM point cloud can achieve better results on background as it can better capture the geometry of the scene while LiDAR point cloud can model moving objects better as it is denser. Our approach combines them to leverage their respective strengths.
Figure 12: Qualitative results of different initial point clouds.
[…]
LiDAR point cloud 31.09 23.31 0.914 0.097
SfM point cloud 31.43 22.80 0.917 0.097
Table 5: Quantitative results of different initial point clouds
Extrapolation results.
In Figure 13, we show some qualitative results of novel view synthesis when the camera is far away from input image sequence on Waymo dataset. Please see our supplementary video for more details.
@@ Appendix B Additional experiments
MARS [55] 19.61 21.06 22.17
Ours 20.29 22.14 22.59
Analysis of optimizing tracked poses.
As discussed in our main paper, we observe that our explicit representation facilitates the optimization of tracked vehicle poses with ease. Herein, we extend our study to explore the impact of an implicit representation on optimizing tracked vehicle poses. The experimental results, as presented in Table 6, indicate that while the inclusion of our pose optimization strategy with implicit representation improves outcomes, there remains a noticeable gap compared to experiments using ground truth tracked poses. However, the proposed method, employing tracked poses from an off-the-shelf tracker, achieves results comparable to those using GT poses. This success can be attributed to the more efficient propagation of gradients through explicit representations in relation to tracked poses.
Figure 13: Qualitative results of novel view synthesis with significant differences from the training set. We translate the camera for 2 meters and rotate it for 30 degrees.

### v2

@@ 4 Implementation details
We train Street Gaussians for 30000 iterations with Adam optimizers [20] following the configurations of 3D Gaussians [19]. The learning rate of translation transformation $\Delta\mathbf{T}_{t}$ and rotation transformation $\Delta\mathbf{R}_{t}$ are set to $5e^{-3}$ and $1e^{-3}$, which decay exponentially to $5e^{-5}$ and $1e^{-5}$ respectively. The resolution of sky cubemap is set to 1024 with learning rate decays from $1e^{-2}$ to $1e^{-4}$ exponentially. All the experiments are conducted on one single RTX 4090 GPU.
We follow [19] to apply adaptive control during optimization. We fix the scale of background model (20 meters in our experiments) and the scale of each object model is determined by the bounding box dimensions. In order to prevent object Gaussians from growing to occluded areas, for each object model we sample a set of points as a probability distribution function. During optimization, Gaussians with sampled points outside the bounding box will be pruned.
@@ 5.1 Experimental Setup
Datasets.  We conduct experiments on Waymo Open Dataset [51] and KITTI benchmarks [15]. The frame rates of both datasets are 10 HZ. On the Waymo Open Dataset, we select 8 recording sequences with large amounts of moving objects, significant ego-car motion and complex lighting conditions. All sequences have a length of around 100 frames. We select every 4th image in the sequence as the test frames and use the remaining for training. As we find that our baseline methods [39, 64] suffer from high memory cost when training with high-resolution images, we downscale the input images to $1066\times 1600$. On KITTI [15] and Vitural KITTI 2 [6], we follow the settings of MARS [64] and evaluate our methods with different train/test split settings. We use the bounding boxes generated by the detector[62] and tracker [63] on Waymo dataset and use the officially provided object tracklets from KITTI.
Baseline methods.  We compare our methods with four recent methods. (1) NSG [39] represents background as multi-plane images and use per-object learned latent codes with a shared decoder to model moving objects. (2) MARS [64] builds the neural scene graph based on Nerfstudio [53]. (3) 3D Gaussians [19] models the scene with a set of anisotropy gaussians. (4) EmerNeRF [69] stratifies scenes into static and dynamic fields, each modeled with a hash grid [36]. Both NSG and MARS are trained and evaluated using ground truth object tracklets. Details of baseline implementations can be found in the supplementary.
Figure 4: Qualitative comparisons results on the Waymo [51] dataset. NSG [39] and MARS [64] often produce blurry and distorted results. 3D GS [19] and EmerNeRF [69] generates ghosting artifacts in regions with moving objects. In contrast, our approach significantly outperforms other methods with high fidelity and sharp details.
3D GS [19] NSG [39] MARS [64] EmerNeRF [69] Ours
@@
PSNR*$\uparrow$ 21.25 24.32 26.54 21.67 30.23
[…]
@@
5.2 Comparisons with the State-of-the-art
Tables 1, 2 present the comparison results of our method with baseline methods [19, 39, 64, 69] in terms of rendering quality and rendering speed. We adopt PSNR, SSIM and LPIPS [77] as metrics to evaluate rendering quality. To better evaluate the rendering quality of moving objects, we project 3D bounding boxes to 2D image plane and calculate the loss only on pixels inside the projected box, which is denoted as PSNR* in our experiments. For all the metrics, our model achieves the best performance among all the methods with a 12.1% increase in PSNR and a 13.9% increase in PSNR*. Moreover, our method renders two magnitudes faster than NeRF-based methods [39, 64, 69]. Although 3D GS is faster than our method, it can only support static scenes and the result of moving objects degrades significantly.
Figure 4 shows the qualitative results of our method and baselines on the Waymo dataset. 3D GS fails to model dynamic objects and EmerNeRF can not generates reasonable results in dynamic regions of novel timestep. Although given ground truth tracking poses, NSG and MARS still suffer from blurry and distorted results due to the lack of capacity of their model when the scene is complex. In contrast, our method can generate high-quality novel views with high fidelity and details.
[…]
Ours w/o LiDAR 34.02 29.53 0.934 0.087
Ours w/o 4DSH 34.36 29.27 0.937 0.081
Ours w/o pose opt. 34.18 28.24 0.935 0.081
Ours w/ GT pose 34.61 29.84 0.937 0.080
Complete model 34.61 30.23 0.938 0.079
Figure 5: Ablation study on tracking pose optimization. The results indicate that optimizing tracked poses improves the quality. “opt.” denotes optimization.
Figure 6: Ablation study on LiDAR point cloud. We show the rendered image and depth of our method with and without LiDAR as input.
@@ 5.3 Ablations and Analysis
We validate our algorithm’s design choices on all selected sequences from the Waymo dataset. Table 3 presents the quantitative results.
Importance of optimizing tracked poses.  Experimental results in Table 3 show that our complete model outperforms the model trained without tracking pose optimization by a large margin, which indicates the effectiveness of our pose optimization strategy. It is interesting to notice that the result of our method is even better than the model trained with ground truth poses, a plausible explanation is that there still exists noise in ground truth annotations.
Visual results of the influence of tracked pose optimization is shown in Figure 5. Treating tracked poses as learnable parameters help the object model synthesize more texture details like the rear of the white vehicle or the logo of the black vehicle and reduce rendering artifacts.
@@
Influence of incorporating LiDAR points.  We evaluate the influence of LiDAR point cloud by comparing our method to a variant with SfM initialization for background and random initialization for moving object as described in Section 3.1. We also disable the LiDAR depth loss in Eq. 9. Table 3 shows that incorporating LiDAR point cloud enhances the results of both background and moving objects. Figure 6 indicates that using LiDAR points helps our model recover more accurate scene geometry and reduce blurry artifacts. It is worth noticing that our method still significantly outperforms baseline methods even without LiDAR input as shown in Tables 3, 4, which proves the efficiency of Street Gaussians under different settings.
Figure 7: Editing operations on the Waymo [51] dataset. Images in the first and second rows represent the results before and after editing. Our method supports various editing operations, including rotation, translation and swapping.
Figure 8: Decomposition results on the Waymo [51] dataset. NSG [39] cannot decompose clean foreground objects while MARS [64] generates floaters in background regions. Instead, our method successfully decomposes the foreground objects and produces high fidelity rendering results.
@@
Figure 9: Visual semantic segmentation results on the KITTI [15] dataset. It can be observed that our method achieves better performance, particularly in ambiguous areas such as shadows, due to our ability to fuse semantic information in 3D.
5.4 Applications
Street Gaussians can be applied to multiple tasks in computer vision including object decomposition, semantic segmentation and scene editing.
Scene editing.  Our instance-aware scene representation enables various types of scene editing operations. We can rotate the heading of the vehicle (Figure 7 (a)), translate the vehicle (Figure 7 (b)) and swap one vehicle in the scene with another one (Figure 7 (c)).
Object Decomposition.  We compare the decomposition results of our method with NSG [39] and MARS [64] under the Waymo dataset. As shown in Figure 8, NSG fails to disentangle foreground objects from the background and the result of MARS is blurry due to the model capacity and lack of regularization. In contrast, our method can produce high-fidelity and clean decomposed results.
Semantic Segmentation.  We compare the quality of our rendered semantic map with the semantic prediction from Video-K-Net [24] on KITTI dataset. Our semantic segmentation model is trained with results from Video K-Net. Qualitative and quantitative results are shown in Figure 9 and Table 4. Our semantic maps achieve better performance thanks to our representation.
6 Conclusion
@@
Acknowledgement.  The authors would like to acknowledge the support from NSFC (No. 623B2091), Li Auto and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
Appendix 0.A More implementation details
0.A.1 Street Gaussians implementations.
Point cloud initialization.  We obtain SfM point cloud of background model by treating camera poses as known parameters and perform point triangulation. As moving objects violate the assumption of multi-view consistency, we ignore these parts by using the mask as shown in Figure 10 during feature extraction. We can directly concatenate SfM and LiDAR point cloud as they are both defined in the world coordinate system.
@@ Appendix 0.A More implementation details
Object semantic.  To merge the one-dimensional scalar $\beta_{o}$ with the M-dimensional vector $\beta_{b}$ of background, we convert $\beta_{o}$ to a M-dimensional one-hot vector for the vehicle label during rendering.
Sky cubemap.  The sky cubemap takes viewing direction $\mathbf{d}$ as input and output sky color $\mathbf{C}_{\text{sky}}$. Let the rendered color and opacity of Gaussians as $\mathbf{C}_{g}$ and $\mathbf{O}_{g}$, the final rendering color $\mathbf{C}$ can be written as:
$\mathbf{C}=\mathbf{C}_{g}+(1-\mathbf{O}_{g})*\mathbf{C}_{\text{sky}}.$ (8)
Loss functions.  As we discussed in the main paper, our total loss function is:
$\mathcal{L}=\mathcal{L}_{\text{color}}+\lambda_{1}\mathcal{L}_{\text{depth}}+\lambda_{2}\mathcal{L}_{\text{sky}}+\lambda_{3}\mathcal{L}_{\text{sem}}+\lambda_{4}\mathcal{L}_{\text{reg}}.$ (9)
  1. 1.
    $\mathcal{L}_{\text{color}}.$ We apply the $\mathcal{L}_{1}$ and D-SSIM loss between rendered and observed images:
    $\mathcal{L}_{\text{color}}=(1-\lambda_{\text{SSIM}})\mathcal{L}_{1}+\lambda_{\text{SSIM}}\mathcal{L}_{\text{D-SSIM}}.$ (10)
    We set $\lambda_{\text{SSIM}}$ to 0.2 following [19].
  2. 2.
    $\mathcal{L}_{\text{depth}}.$ We apply the $\mathcal{L}_{1}$ loss between rendered depth $\mathbf{D}$ and the LiDAR measurement’s depth $\mathbf{D}^{\text{lidar}}$:
    $\mathcal{L}_{\text{depth}}=\sum||\mathbf{D}-\mathbf{D}^{\text{lidar}}||_{1}$ (11)
    We optimize 95% of the pixels with smallest depth error to prevent noisy LiDAR observations from affecting the optimization [71]. $\lambda_{1}$ is set to 0.01.
  3. 3.
    $\mathcal{L}_{\text{sky}}.$ We apply the binary cross entropy loss between rendered opacity $\mathbf{O}_{g}$ and predicted sky mask $\mathbf{M}_{\text{sky}}$:
    $\mathcal{L}_{\text{sky}}=-\sum((1-\mathbf{M}_{\text{sky}})\text{log}\mathbf{O}_{g}+\mathbf{M}_{\text{sky}}\text{log}(1-\mathbf{O}_{g}))$ (12)
    $\mathbf{M}_{\text{sky}}$ is generated by Grounded SAM [44]. To be specific, we first get 2D boxes by entering text "sky" to Grounding Dino [31]. Then we input the boxes as prompt to SAM [21] and obtain the predicted sky mask. $\lambda_{2}$ is set to 0.05.
  4. 4.
    $\mathcal{L}_{\text{sem}}.$ We apply the per-pixel softmax-cross-entropy loss between rendered semantic logits and predicted 2D semantic segmentation [24]. In order to prevent noisy input semantic labels from influencing the scene geometry [49], we only perform backpropagation to semantic logits $\beta$ for $\mathcal{L}_{\text{sem}}$. $\lambda_{3}$ is set to 0.1.
  5. 5.
    $\mathcal{L}_{\text{reg}}.$ The regularization term in our loss function is defined as an entropy loss on the accumulated alpha values of decomposed foreground objects $\mathbf{O}_{\text{obj}}$:
    $\mathcal{L}_{\text{reg}}=-\sum(\mathbf{O}_{\text{obj}}\text{log}\mathbf{O}_{\text{obj}}+(1-\mathbf{O}_{\text{obj}})\text{log}(1-\mathbf{O}_{\text{obj}}))$ (13)
    We add this loss after the adaptive control process to help our model better distinguishes foreground and background. Figure 11 shows the qualitative results, which demonstrates the effect of this regularization term. $\lambda_{4}$ is set to 0.1.
Hyperparameters.  In practice, we set the number of fourier coefficients $k$ as 5 to maintain a balance between performance and storage cost. Due to the relatively less intense view-dependent effect on urban scene compared to dataset in 3D Gaussian [19], we reduce the SH degree to 1 to prevent overfitting. We set the voxel size to 0.15m when performing downsampling for LiDAR point cloud.
0.A.2 Baselines implementations.
We give detailed descriptions of our baseline implementations.
  1. 1.
    Neural Scene Graph [39]. We use the official implementation and try a variant where each moving object is modeled by a separate NeRF network instead of a shared decoder. The best result is reported for each scene.
  2. 2.
    Mars [64]. We use the official implementation and try a variant where each moving object is modeled by a separate Nerfacto [53] model. We choose the appearance embedding of nearest training frame as input to each test frame. The best result is reported for each scene.
  3. 3.
    3D Gaussians [19]. We initialize the point cloud by running Colmap [46] with known camera parameters from the dataset. We find that the number of SfM point cloud generated by Colmap is usually less than 1K when the ego-vehicle has little motion. We use LiDAR points to initialize 3D Gaussians for these cases to get reasonable results.
  4. 4.
    EmerNeRF [69]. We run the official code under the setting incorporating dynamic encoder, flow encoder and feature lifting. We use the same sky mask obtained from Grounded SAM [44] for fair comparison.
0.A.3 Evaluations
Figure 10 visually illustrates the calculation method of the PSNR* metric in our experiments. We expand each bounding box by 1.5 times in both length and width dimensions to ensure it fully covers the object. For fair comparison, both our method and the baselines are evaluated using the mask obtained from object tracklets provided by the dataset.
@@ 0.A.3 Evaluations
Appendix 0.B Additional experiments
Analysis of optimizing tracked poses.  As discussed in our main paper, we observe that our explicit representation facilitates the optimization of tracked vehicle poses with ease. Herein, we extend our study to explore the impact of an implicit representation [64] on optimizing tracked vehicle poses. The experimental results, as presented in Table 5, indicate that while the inclusion of our pose optimization strategy with implicit representation improves outcomes, there remains a noticeable gap compared to experiments using ground truth tracked poses. However, the proposed method, employing tracked poses from an off-the-shelf tracker, achieves results comparable to those using GT poses. This success can be attributed to the more efficient propagation of gradients through explicit representations in relation to tracked poses.
@@ Appendix 0.B Additional experiments
MARS [64] 24.38 25.83 26.94
Ours 25.86 27.98 28.02
Analysis of point cloud initialization.  We have demonstrated the importance of including LiDAR point cloud during initialization. However, the LiDAR points can not cover the entire scene, especially for far away regions. As a result, the SfM point cloud is also crucial for the reconstruction of dynamic urban scene. We perform an ablation study by using only LiDAR point cloud to initialize the background model. As shown in Figure 12, although LiDAR points can help recover texture details in near regions like the road, it cannot restore some areas not covered by LiDAR points, such as trees on the other side of the road or distant road signs. The rendering results without SfM points is even worse than the one without LiDAR points as illustrated in Table 6. Our approach combines them as input to leverage their respective strengths.
Analysis of sky modeling.  As shown in Figure 13 and Table 6, using a separate cube map to model the sky can help better recover detail areas, while avoiding some foreground objects being obscured by the gaussians representing sky regions.
Table 6: More ablation studies on point cloud initialization and sky modeling. We show the quantitative results on two scenes from Waymo dataset with large scale background and many thin structures.
[…]
Ours 32.63 0.928 0.083
Ours w/o LiDAR 30.72 0.920 0.100
Ours w/o SfM 29.97 0.911 0.106
Ours w/o Sky modeling 31.12 0.921 0.100
Figure 12: Effect of incorporating SfM points on novel view synthesis results.
Figure 13: Effect of modeling sky with cubemap on novel view synthesis results.
Extrapolation results.  In Figure 14, we show some qualitative results of novel view synthesis under the setting of lane changes on Waymo dataset. Our method can produce high-quality results although the rendering viewpoint is far away from input sequence.
Figure 14: Qualitative results of novel view synthesis with significant differences from the input frames. In each scene we shift the camera by 2 meters.
Qualitative results on the KITTI dataset.  In Figure 15, we show the comparison results with NSG [39] and MARS [64] on the KITTI [15] dataset.
Figure 15: Qualitative comparison results on the KITTI [15] dataset.
Decomposition results on the KITTI dataset.  In Figure 16, we show the qualitative comparisons of decomposition with NSG [39] and Panoptic Neural Fields [23] on the KITTI [15] dataset.
Figure 16: Decomposition results on the KITTI [15] dataset.
Appendix 0.C Limitations
Street Gaussians also has some known limitations. 1) Our method is limited to reconstructing rigid dynamic scenes, such as static streets with only moving vehicles, and cannot handle non-rigid dynamic objects like walking pedestrians. Future work could consider employing more complex dynamic scene modeling methods [73], to address this issue. 2) the proposed method is dependent on the recall rate of off-the-shelf trackers. If some vehicles are missed, our pose optimization strategy cannot compensate for this. Obtaining continuous tracklets through methods like 2D tracking can alleviate the problem and modeling dynamic urban scenes without object tracklets remains an interesting problem. 3) Street Gaussians still requires per-scene optimization. We consider predicting generalizable 3D Gaussians in feed-forward manner as a future work.